OpenStudy (anonymous):

HELP ASAP Find the average rate of change for f(x) from x=4 to x=7 and my equation is f(x)=2(x-10)^2+15 please I really need help

OpenStudy (anonymous):

you are being asked to compute \[\frac{f(7)-f(4)}{7-4}\]

OpenStudy (anonymous):

After that what do I do? And thank you so much

OpenStudy (anonymous):

after that you go have a beer or soft drink if you are under age that is the answer

OpenStudy (anonymous):

i mean you do need to compute those numbers and divide what is \(f(7)\) ?

OpenStudy (anonymous):

I know the formula but I don't know how to plug in the numbers

OpenStudy (anonymous):

\[f(x)=2(x-10)^2+15\] \[f(7)=2(7-10)^2+15\] what do you get?